# On the PM2.5 – Mortality Relationship: A Bayesian Dynamic Model for Spatio-Temporal Confounding

### v1

@@ 3 The Proposed Model
If $f(\cdot;\bm{\gamma}_{f})$ is a constant function, i.e. $f(X_{it};\bm{\gamma}_{f})=\gamma_{f0}$, then $\beta_{1t}=0$ and $\beta_{0it}=\beta_{0i}+\gamma_{f0}+Z_{it}$.
@@
If $f(\cdot;\bm{\gamma}_{f})$ is a linear function, i.e. $f(X_{it};\bm{\gamma}_{f})=\gamma_{f0}+\gamma_{f1}X_{it}$, then $\beta_{1t}=\gamma_{f1}$ and $\beta_{0it}=\beta_{0i}+\gamma_{f0}+\gamma_{f1}\widehat{X_{t}}+Z_{it}$. The model proposed by Janes et al. (2007) can be cast in this case, with $\gamma_{f0}=0$. The authors, however, assume further that the intercepts are constant, so the overall trend in the exposure is included as a covariate. They also include a smooth function of time in their specification to control for seasonality and long-term trends common to all areas.
3.1 Specification of the Time-Varying Coefficients
@@
We implement a DGLM with random walk processes to model the temporal evolution of the regression coefficients $\beta_{0it}$ and $\beta_{1t}$. We propose the following state-space formulation for the latent process and the regression coefficients, for $t=1,\dots,{T}$:
$\bm{\vartheta}_{t}$ $=\mathbf{o}_{t}+\mathbf{\tilde{X}}_{t}\tilde{\bm{\beta}}_{t}+{{}^{m}\mathbf{Z}}_{t}\bm{\alpha}_{m}+\tilde{{\bf u}}_{t}\,,\quad\tilde{{\bf u}}_{t}\sim\mathcal{N}({\bf 0},\tilde{\tau}_{u}^{2}{\bf I}_{N})$ (7)
$\tilde{\bm{\beta}}_{t}$ $=\tilde{\bm{\beta}}_{t-1}+\mathbf{w}_{t},\quad{\bf w}_{t}\sim\mathcal{N}({\bf 0},\bm{\varSigma}_{w})\,,$ (8)
where $\bm{\vartheta}_{t}=(\vartheta_{1t},\dots,\vartheta_{{N}t})^{\prime}$, $\mathbf{o}_{t}=(o_{1t},\dots,o_{{N}t})^{\prime}$,
${}^{m}\mathbf{Z}_{t}=\begin{bmatrix}{}^{m}\mathbf{Z}_{1t}^{\prime}\\ \vdots\\ {}^{m}\mathbf{Z}_{{N}t}^{\prime}\end{bmatrix}\,,$
and $\tilde{\bm{\beta}}_{t}=\left(\beta_{01t},\dots,\beta_{0{N}t},\beta_{1t}\right)^{\prime}$ is an $(N+1)$-dimensional vector. The observation matrix is $\mathbf{\tilde{X}}_{t}=\begin{bmatrix}\mathbf{I}_{N}&\mathbf{\tilde{X}}_{\,\bm{\cdot}t}\end{bmatrix}$, where
@@
$\mathbf{\tilde{X}}_{\,\bm{\cdot}t}=\begin{bmatrix}X_{1t}-\widehat{X_{t}}\\ \vdots\\ X_{{N}t}-\widehat{X_{t}}\end{bmatrix}\,.$
4 Bayesian Inference and Computations
@@
The first stages of the Bayesian hierarchical model are defined by Equations (1) and (7)–(8). To complete it, a prior distribution must be assigned on the parameter set, $\bm{\Theta}=\{\tilde{\bm{\beta}}_{0},\bm{\alpha}_{m},\bm{\varSigma}_{w},\tilde{\tau}_{u}^{2}\}$.
@@ 4.1 Prior Specification
Independent prior distributions are imposed on each parameter in $\bm{\Theta}$. As for the vector of time-varying coefficients, $\tilde{\bm{\beta}}_{t}$, distributional assumptions are required on the initial state (i.e., when $t=0$), and on the innovation error covariance matrix, $\bm{\varSigma}_{w}$. We impose that $\tilde{\bm{\beta}}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{V}_{\beta})$, where $\mathbf{V}_{\beta}=\begin{bmatrix}10\mathbf{I}_{N}&\mathbf{0}\\ \mathbf{0}&4\times 10^{-6}\end{bmatrix}$.
The prior specification for $\bm{\varSigma}_{w}$ requires more care. We assume that this matrix is block-diagonal, i.e. $\bm{\varSigma}_{w}=\begin{bmatrix}\bm{\varSigma}_{w0}&\mathbf{0}\\ \mathbf{0}&\sigma_{w1}^{2}\end{bmatrix}$. Given that the area-specific intercepts should capture the effect of the unmeasured confounder, $Z_{it}$, we expect them to be spatially correlated, so $\bm{\varSigma}_{w0}$ is an ${N}\times{N}$ matrix for which an inverse-Wishart prior is assigned, i.e. $\bm{\varSigma}_{w0}\sim\mathcal{IW}(a_{w0},\mathbf{S}_{w0})$, with $a_{w0}={N}+2$ and $\mathbf{S}_{w0}=k_{0}\mathbf{I}_{N}$. This specification allows for a free-form and data-driven covariance structure, and facilitates inference, since the posterior full conditional distribution of $\bm{\varSigma}_{w0}$ is available in closed-form (Brown et al., 1994). An inverse-gamma prior is imposed on $\sigma_{w1}^{2}$, i.e. $\sigma_{w1}^{2}\sim\mathcal{IG}(a_{w1},b_{w1})$. Other specifications were also considered (e.g., $\bm{\varSigma}_{w}$ was assumed to be diagonal), but they did not provide good mixing of the posterior samples chains. Recall that capturing gradual changes is a desired feature of the proposed model, and the matrix $\bm{\varSigma}_{w}$ plays a crucial role to this end. For instance, if $k_{0}$ is “large”, the short-term variation in the intercepts becomes unlikely pronounced, whereas if $k_{0}$ is “small”, the “true” variation is captured only in part. The definition of “large” and “small” mainly depends on the sample size and on how much the prior dominates the likelihood. Note also that extreme values of $k_{0}$ may lead to numerical non-positive definiteness of $\bm{\varSigma}_{w0}$.
Therefore, several values for the hyperparameters $k_{0}$, $a_{w1}$ and $b_{w1}$ are considered to evaluate the sensitivity of the posterior distribution to the prior. In the case of both simulated and real data (see Sections 5 and 6, respectively), we find that $k_{0}=10^{-3}$, $a_{w1}=11$, and $b_{w1}=10^{-6}$ provide a good balance between the prior and the likelihood.
Finally, non-informative normal priors are imposed on the coefficients of the measured confounders, $\bm{\alpha}_{m}\sim\mathcal{N}(\mathbf{0},10^{6}\mathbf{I}_{{p_{m}}})$. Following Wakefield (2007), it is assumed that the residual relative risks $\tilde{u}_{it}$, follow a log-t distribution with two degrees of freedom, and with $95\%$ of prior mass within the interval $(0.98,1.02)$, hence we obtain the prior $\tilde{\tau}_{u}^{2}\sim\mathcal{IG}(1,2.1\times 10^{-5})$.
@@ 4.2 Computations
As first step of the analysis, the overall temporal trend, $\widehat{X_{t}}$, is obtained as the fitted values from the model in Equation (4), where $h(t;df)$ represents a natural cubic spline of time with $df$ degrees of freedom. This two-step strategy should be preferred to a joint modeling of the exposure and the outcome, as it allows avoiding feedback from the outcome to the exposure that may lead to biased estimators of the exposure effect (see, e.g., Stephens et al., 2023).
An MCMC scheme is then implemented to simulate from the posterior distribution of $\bm{\Theta}$. The algorithm proposed by Chan and Jeliazkov (2009) is used to sample from the full conditionals of the time-varying coefficients. It is an efficient simulation algorithm that allows to sample all the regression coefficients (for all the time instants) in a single step. The remaining full conditional distributions are composed of inverse-gamma, and inverse-Wishart distributions. The latent process $\vartheta_{it}$ is an exception and is sampled using a Metropolis-Hastings step. Further details can be found in Section D of the Supplementary Material.
@@ 5 Analysis of Artificial Data
To validate the proposed approach, we conduct two simulation studies, both inspired by the real data example presented in Section 2. In the first one, which is described next, we want to assess the ability of the proposed model to mitigate the confounding bias.
@@
We consider ${N}=100$ contiguous HDs of Piemonte and Lombardia, then we assume that ${T}=300$, that there are no measured confounders (i.e., ${p_{m}}=0$) and that the exposure $X_{it}$ and the unmeasured confounder $Z_{it}$ are correlated vector autoregressive processes of order $1$ (VAR($1$)) with a linear correlation coefficient $\rho_{xz}={\rm\mathbb{C}or}(X_{it},Z_{it})=0.5$, for $i=1,\dots,{N}$, and $t=1,\dots,{T}$, and with spatially-structured, temporally independent innovations (Cressie and Wikle, 2015, Sahu, 2022). A proper conditional autoregressive (PCAR) specification (Banerjee et al., 2014, Besag, 1974, Sahu, 2022) is assumed to model the spatial dependence of these innovations. The marginal spatio-temporal means of exposure and unmeasured confounder are generated from Fourier basis expansions with period ${T}$. The correlation matrix of $X_{it}$ is parameterized by spatial and temporal autocorrelation parameters, i.e. $\phi_{x}^{S}\in(0,1)$ and $\phi_{x}^{T}\in(0,1)$, respectively. Similarly, $\phi_{z}^{S}$ and $\phi_{z}^{T}$ represent the autocorrelation parameters of $Z_{it}$, which is sampled conditional on $X_{it}$. These four parameters represent the strength of the spatial and temporal dependence, and define three scenarios that are summarized in Table 1. The outcome variable is sampled from the Poisson distribution in Equation (1) using the following linear predictor:
$\vartheta_{it}=o_{it}+\beta_{0i}+\beta_{1t}X_{it}+Z_{it}+u_{it},\quad u_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{u}^{2})\,,$
where $\tau_{u}^{2}=10^{-5}$. Additional information on the data generation mechanism is provided in Section E of the Supplementary Material.
@@
The proposed approach is compared to the Null model and to the GLMadj model, where natural cubic spline of time with $7$ degrees of freedom is used for the latter. Following Peng et al. (2006), the degrees of freedom for the spline are chosen by fitting several models with different numbers of bases and selecting the minimum number from which the estimates appear stabilized.
@@
Figure 5 depicts the results in terms of percent change of risk associated with a 10-unit increase in the exposure, defined as $P_{t}=100(\exp\{10\beta_{1t}\}-1)$. For each scenario, the black solid line represents the case where true exposure effect is a constant (top), a periodic (middle), or a more complex function of time (bottom). The estimate of $P_{t}$ obtained from the proposed DGLM is shown in red, with the solid line indicating its average, and the shaded area indicating the percentiles $2.5$ and $97.5$ of its distribution over 100 replicates.
@@
The dashed gray and dashed green lines represent the estimated percent changes of risk from the Null and GLMadj models, averaged over the 100 replicates. Whereas the Null model is evidently biased in all cases, the GLMadj model and the proposed approach significantly reduce the confounding bias in Scenario S1, although without entirely removing it. In Scenario S2, the confounding bias reduction is still achieved by the DGLM compared to the Null model, but to a lesser extent. This could be because the unmeasured confounder varies at high frequencies in time, so it cannot be captured by the time-varying intercepts that only account for gradual changes. Also, the difference $X_{it}-\widehat{X_{t}}$ is approximately as rough as $Z_{it}$, and, due to their correlation, a portion of the variability of the unmeasured confounder is explained by the local exposure effect. Scenario S3 is very similar to Scenario S1, suggesting that the lack of smoothness in space of the unmeasured confounder does not seem to alter the performance (in terms of confounding bias) of the proposed approach.
@@
Figure 5: Simulation Study 1: results are expressed in terms of percent change of risk associated with a 10-unit increase in the exposure, $P_{t}$. The black solid line represents the true exposure effect (whose shape differs in each row), while the dashed gray line represents the estimated exposure effect from the Null model. The red solid line and the shaded area represent the average and the percentiles $2.5$ and $97.5$ of the posterior means of $P_{t}$ obtained from the proposed DGLM over 100 replicated datasets.
These results are confirmed also by Table E.4 in the Supplementary Material, where the average mean squared error ($MSE_{avg}$), the maximum absolute bias ($MAB$), and the average mean absolute error ($MAE_{avg}$) are reported for each model and each scenario (see the Supplementary Material, Section E.2, for definitions). The GLMadj model has the lowest values when the exposure effect is constant. This is likely because the GLMadj model correctly assumes a constant exposure–outcome relationship, whereas the proposed model learns its shape from the data. Nevertheless, in the top subplots of Figure 5, horizontal straight lines can be drawn through the red shaded areas, indicating that the DGLM correctly identifies a constant exposure effect. In the remaining cases, the proposed model seems to perform best, as it is able to capture the temporal variations of the true exposure effect.
Model Reparameterization
Section E.3 of the Supplementary Material also motivates, from a practical viewpoint, the decomposition of the exposure effect into the two components described in Equation (3). We compare the proposed DGLM with a reparameterized model that does not remove the overall trend from the exposure. The results suggest that the proposed approach is more effective in reducing the confounding bias and almost always has lower $MSE_{avg}$, $MAB$ and $MAE_{avg}$ values.
@@
Simulation Study 2: Recovery of the Local Exposure Effect ($f^{\prime}$)
@@
In the second simulation study, described in Section F of the Supplementary Material, we want to assess whether the proposed DGLM is able to recover $f^{\prime}(\cdot;\bm{\gamma}_{f})$, i.e. the first derivative of the unknown functional form. The data are generated using Equations (1)–(2), and we show that the DGLM is able to capture the dynamics of the exposure effect, but there remains a residual bias which is only due to confounding (i.e., not due to Taylor’s approximation).

### v2

@@ 3 The Proposed Model
If $f(\cdot;\bm{\gamma}_{f})$ is a constant function, i.e. $f(X_{it};\bm{\gamma}_{f})=\gamma_{f0}$, then $\beta_{1it}=0$ and $\beta_{0it}=\beta_{0}+\gamma_{f0}+Z_{it}$.
@@
If $f(\cdot;\bm{\gamma}_{f})$ is a linear function, i.e. $f(X_{it};\bm{\gamma}_{f})=\gamma_{f0}+\gamma_{f1}X_{it}$, then $\beta_{1it}=\gamma_{f1}$ and $\beta_{0it}=\beta_{0}+\gamma_{f0}+\gamma_{f1}\widehat{X}_{it}+Z_{it}$. The model proposed by Greven et al. (2011) can be cast in this case, with $\gamma_{f0}=0$. The authors, however, assume further that the intercepts do not depend on time.
3.1 Specification of the Spacetime-Varying Coefficients
Although it is possible to model $\beta_{0it}$ and $\beta_{1it}$ directly, we decompose these spacetime-varying coefficients as follows to enhance prior elicitation and interpretability (see also Bakar et al., 2015, Cai et al., 2013, Paez et al., 2008):
$\beta_{kit}=\overline{\delta}_{k}+\delta_{ki}+\delta_{kt}+\delta_{kit}^{*}\,,\quad k=0,1\,,$
where $\overline{\delta}_{k}$ is the global effect of the $k$th covariate over the whole study period, $\delta_{ki}$ is the area-specific effect of the $k$th covariate, $\delta_{kt}$ is the time-specific effect of the $k$th covariate, and $\delta_{kit}^{*}$ is the spatio-temporal interaction effect of the $k$th covariate. For each $i$, $\delta_{ki}$ can be interpreted as the average (across time) spatial deviation from the global effect. In particular, $\delta_{1i}$ captures time-independent spatial heterogeneity in the exposure–outcome association at the local scale of variation (recall that the regressor is $X_{it}-\widehat{X}_{it}$). It can inform about where populations are most vulnerable to the effects of air pollution.
For each time instant, $\delta_{kt}$ can be interpreted as the average (across space) temporal changes from the global effect. Hence, $\delta_{1t}$ represents the temporal evolution of the association of interest, and highlights when peaks can occur. The interaction effect, $\delta_{kit}^{*}$, captures the area- and time-specific deviations from the global effect that are left after accounting for the temporal and spatial effects. The simulation scenarios in Section 6 were built based on possible values of $\beta_{1it}$ that are very small in real applications. Hence, given the tiny contribution of the term $\beta_{1it}(X_{it}-\widehat{X}_{it})$ to the augmented-data likelihood (see Section 4.2), it is difficult to recover the interaction effect of the exposure, even in the absence of confounding. Therefore, we assume that $\delta_{1it}^{*}$ is zero. This assumption is in line with previous studies (see, e.g., Bakar et al., 2015, Paez et al., 2008).
@@ 3.1 Specification of the Spacetime-Varying Coefficients
In SDGLMC, we implement a DGLM with random walk (RW) processes to model the temporal effects, $\delta_{0t}$ and $\delta_{1t}$, and the interaction effect, $\delta_{0it}^{*}$. We propose the following state-space formulation for the latent process and the regression coefficients, for $t=1,\dots,{T}$:
$\bm{\vartheta}_{t}$ $=\mathbf{o}_{t}+{\bf 1}_{N}(\overline{\delta}_{0}+\delta_{0t})+\tilde{\bm{\delta}}_{0}+\bm{\delta}_{0t}^{*}+\mathbf{\tilde{X}}_{t}(\overline{\delta}_{1}+\delta_{1t})+\mathbf{\tilde{X}}_{t}\odot\tilde{\bm{\delta}}_{1}+{\mathbf{M}}_{t}\bm{\alpha}+\tilde{{\bf u}}_{t}$ (7)
$\delta_{0t}$ $=\delta_{0,t-1}+w_{0t},\quad w_{0t}\sim\mathcal{N}(0,\sigma_{w0}^{2})$ (8)
$\bm{\delta}_{0t}^{*}$ $=\bm{\delta}_{0,t-1}^{*}+\mathbf{w}_{0t}^{*},\quad\mathbf{w}_{0t}^{*}\sim ICAR(\sigma_{w0}^{*2})$ (9)
$\delta_{1t}$ $=\delta_{1,t-1}+w_{1t},\quad w_{1t}\sim\mathcal{N}(0,\sigma_{w1}^{2})$ (10)
where the symbol $\odot$ denotes the Hadamard product, $\bm{\vartheta}_{t}=(\vartheta_{1t},\dots,\vartheta_{{N}t})^{\prime}$, $\mathbf{o}_{t}=(o_{1t},\dots,o_{{N}t})^{\prime}$, $\tilde{\bm{\delta}}_{k}=(\delta_{k1},\dots,\delta_{k{N}})^{\prime}$ are the spatial-specific effects for $k=0,1$, $\tilde{{\bf u}}_{t}=(\tilde{u}_{1t},\dots,\tilde{u}_{{N}t})^{\prime}\sim\mathcal{N}({\bf 0},\tilde{\tau}_{u}^{2}{\bf I}_{N})$, $\bm{\delta}_{0t}^{*}=(\delta_{01t}^{*},\dots,\delta_{0{N}t}^{*})^{\prime}$, and the observation matrices are as follows:
@@
$\mathbf{\tilde{X}}_{t}=\begin{bmatrix}X_{1t}-\widehat{X}_{1t}\\ \vdots\\ X_{{N}t}-\widehat{X}_{{N}t}\end{bmatrix}\quad\text{and}\quad\mathbf{M}_{t}=\begin{bmatrix}\mathbf{M}_{1t}^{\prime}\\ \vdots\\ \mathbf{M}_{{N}t}^{\prime}\end{bmatrix}\,.$
Note that (9) is one of five alternatives that will be discussed further in Section 7. Similarly to $\mathbf{w}_{0t}^{*}$, to capture the spatial correlation among the area-specific effects, $\tilde{\bm{\delta}}_{0}$ and $\tilde{\bm{\delta}}_{1}$ are modeled as intrinsic conditional autoregressive (ICAR; see, e.g., Banerjee et al., 2014) processes, denoted as $ICAR(\sigma_{\delta k}^{2})$, with density proportional to
$p(\tilde{\bm{\delta}}_{k})\propto\exp\biggl{\{}-\frac{1}{2\sigma_{\delta k}^{2}}\tilde{\bm{\delta}}_{k}^{\prime}(\mathbf{D}-\mathbf{W})\tilde{\bm{\delta}}_{k}\biggr{\}},\quad k=0,1\,,$
where $\mathbf{W}$ is the spatial adjacency matrix with elements $(\mathbf{W})_{ij}=1$ if areal units $i$ and $j$ are neighbors, and 0 otherwise, $\mathbf{D}$ is a diagonal matrix with the sum of the rows of $\mathbf{W}$, and $\sigma_{\delta k}^{2}$ is a variance parameter. Such structure has been previously considered by Franco-Villoria et al. (2019, 2022) to model spatially varying coefficients.
4 Bayesian Inference and Computations in SDGLMC
@@
The first stages of the Bayesian hierarchical model are defined by Equations (1) and (7)–(10). To complete it, a prior distribution must be assigned on the parameter set, $\bm{\Theta}=\{\overline{\delta}_{0},\delta_{00},\bm{\delta}_{00}^{*},\overline{\delta}_{1},\delta_{10},\bm{\alpha},\sigma_{w0}^{2},\sigma_{\delta 0}^{2},\sigma_{w0}^{*2},\sigma_{w1}^{2},\sigma_{\delta 1}^{2},\tilde{\tau}_{u}^{2}\}$.
@@ 4.1 Prior Specification
Independent prior distributions are imposed on each parameter in $\bm{\Theta}$. As for the time-varying coefficients, distributional assumptions are required on the initial states (i.e., when $t=0$), and on the variances of innovation errors. We impose that $\delta_{k0}\sim\mathcal{N}(0,V_{\delta k})$, $k=0,1$, where $V_{\delta 0}=4$ and $V_{\delta 1}=4\times 10^{-6}$. Also, $\bm{\delta}_{00}^{*}\sim\mathcal{N}(\mathbf{0},\mathbf{V}_{\delta}^{*})$, with $\mathbf{V}_{\delta}^{*}=10\mathbf{I}_{N}$.
Inverse-gamma priors are imposed on $\sigma_{w0}^{*2}\sim\mathcal{IG}(a_{w0}^{*},b_{w0}^{*})$, and similarly on $\sigma_{wk}^{2}\sim\mathcal{IG}(a_{wk},b_{wk})$ and $\sigma_{\delta k}^{2}\sim\mathcal{IG}(a_{\delta k},b_{\delta k})$, for $k=0,1$. For $k=0$, all the hyperparameters are assumed to be $0.01$. For $k=1$, choosing the hyperparameters requires more care. Recall that capturing gradual changes is a desired feature of the proposed model, and the variances $\sigma_{wk}^{2}$ and $\sigma_{\delta 1}^{2}$ play a crucial role to this end. For instance, if their prior means are “large”, the variations in the temporal and spatial effects, respectively, become unlikely pronounced, whereas if their prior means are “small”, the “true” variation is captured only in part. The definition of “large” and “small” mainly depends on the sample size and on how much the prior dominates the likelihood.
Therefore, several values for the hyperparameters in both inverse-gamma distributions are considered to evaluate the sensitivity of the posterior distribution to the prior. In the case of both simulated and real data (see Sections 6 and 7, respectively), we find that $a_{w1}=10$, $b_{w1}=10^{-3}$, and $a_{\delta 1}=1,b_{\delta 1}=0.01$ provide a good balance between the prior and the likelihood.
Finally, the global effects, $\overline{\delta}_{0}$ and $\overline{\delta}_{1}$, and the coefficients of the measured confounders, $\bm{\alpha}$, are assumed independently and normally distributed with mean zero and variance $10^{6}$. Following Wakefield (2007), it is assumed that the residual relative risks, $\tilde{u}_{it}$, follow a log-t distribution with two degrees of freedom, and with $95\%$ of prior mass within the interval $(0.98,1.02)$, hence we obtain the prior $\tilde{\tau}_{u}^{2}\sim\mathcal{IG}(1,2.1\times 10^{-5})$.
@@ 4.2 Computations
As a first step of the analysis, the overall temporal trend, $\widehat{X}_{it}$, is obtained as the fitted values from the model in Equation (4), where $h({\bf s}_{i},t;df)$ represents a principal spline of space and time (see, e.g., Mardia et al., 1998, Sahu and Mardia, 2005, Fontanella et al., 2013) with $df$ bases. This two-step strategy should be preferred to a joint modeling of the exposure and the outcome, as it allows avoiding feedback from the outcome to the exposure that may lead to biased estimators of the exposure effect (see, e.g., Stephens et al., 2023).
An MCMC scheme is then implemented to simulate from the posterior distribution of $\bm{\Theta}$. The algorithm proposed by Chan and Jeliazkov (2009) is used to sample from the full conditionals of the time-varying coefficients. It is an efficient simulation algorithm that allows to sample all the regression coefficients (for all the time instants) in a single step. The remaining full conditional distributions are inverse-gamma distributions. The latent process $\vartheta_{it}$ is an exception and is sampled using Metropolis-Hastings steps. To avoid identifiability issues, all time-specific, area-specific and interaction effects are centered at zero during each MCMC iteration (see, e.g., Knorr-Held, 2000, Franco-Villoria et al., 2022). Further details can be found in Section C of the Supplementary Material.
5 Competing Approaches
In this section, we briefly describe the competing approaches that are considered in the simulation study (in Section 6) and in the real data application (in Section 7). Similarly to SDGLMC, they are all based on the assumption that the outcome follows a Poisson distribution. The main difference lies in how the exposure effect is modeled, and how unmeasured confounding is accounted for.
The linear predictor is a function of exposure and measured confounders. To control for unmeasured confounding, additional covariates and basis functions are included in the linear predictor (see, e.g., Peng et al., 2006, Peng and Dominici, 2008). Several approaches from the literature are then fitted under the Bayesian paradigm: (i) the Null model, which does not account for unmeasured confounding; (ii) the GLMadj model, which adjusts for unmeasured confounding through a natural cubic spline of time (see, e.g., Peng and Dominici, 2008, Zhou et al., 2021, Masselot et al., 2022); (iii) the JDZ model by Janes et al. (2007), which is similar to the previous approach but further decomposes the exposure effect in two orthogonal parts; (iv) the Dummy model (see, e.g, Dai et al., 2014, Klompmaker et al., 2023), which allows for season-specific exposure effects; and (v) the Periodic model by Peng et al. (2005) and Chen et al. (2013), which assumes that the exposure effect is a linear combination of trigonometric functions. A hierarchical structure is assumed for each of them, with spatial-specific coefficient vectors being independently and identically distributed. Further details are provided in Section D of the Supplementary Material.
@@ 6 Analysis of Artificial Data
To validate the proposed approach, we conduct two simulation studies, both inspired by the real data example presented in Section 2. In the first one, which is described next, we want to assess the ability of SDGLMC to mitigate the confounding bias.
@@
We consider the ${N}=117$ contiguous HDs of Piemonte and Lombardia, then we assume that ${T}=600$ and there are no measured confounders (i.e., ${p_{m}}=0$). The exposure $X_{it}$ and the unmeasured confounder $Z_{it}$ are correlated vector autoregressive processes of order $1$ (VAR($1$)) with a linear correlation coefficient of 0.5, and with spatially-structured, temporally independent innovations (Cressie and Wikle, 2015, Sahu, 2022). A proper conditional autoregressive (PCAR) specification (Banerjee et al., 2014, Besag, 1974, Sahu, 2022) is assumed to model the spatial dependence of these innovations. The marginal spatio-temporal means of exposure and unmeasured confounder are generated from Fourier basis expansions. The correlation matrix of $X_{it}$ is parameterized by spatial and temporal autocorrelation parameters, i.e. $\phi_{x}^{S}\in(0,1)$ and $\phi_{x}^{T}\in(0,1)$, respectively. Similarly, $\phi_{z}^{S}$ and $\phi_{z}^{T}$ represent the autocorrelation parameters of $Z_{it}$, which is sampled conditional on $X_{it}$. These four parameters represent the strength of the spatial and temporal dependence, and define three scenarios that are summarized in Table 1.
It is worth noting that, in the concurvity setting, $X_{it}$ and $Z_{it}$ are considered as fixed unknown functions, particularly when exposure and unmeasured confounder vary at large scales. The stochastic generative model we use continues to be valuable within this framework of fixed functions because the realizations of $X_{it}$ and $Z_{it}$ provide reasonable values for exposure and unmeasured confounder that could arise in real-world applications where a reasonable stochastic process is absent (Paciorek, 2010).
The outcome variable is sampled from the Poisson distribution in Equation (1) using the following linear predictor:
${\color[rgb]{0,0,0}\vartheta_{it}=o_{it}+\beta_{0}+\beta_{1it}X_{it}+Z_{it}+u_{it},\quad u_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{u}^{2})\,,}$
where $\tau_{u}^{2}=10^{-3}$. Additional information on the data generation mechanism is provided in Section E of the Supplementary Material.
@@
The proposed approach is compared to the Null model and to the GLMadj model. The results from the other competing approaches are not reported here, but they can be found in Section E of the Supplementary Material. A natural cubic spline of time with $6$ degrees of freedom is used to account for unmeasured confounding in GLMadj. Following Peng et al. (2006), the degrees of freedom for the spline are chosen by fitting several models with different numbers of bases and selecting the minimum number from which the estimates appear stabilized.
@@
Figure 3 depicts the results in terms of percent change of risk associated with a 10-unit increase in the exposure, specifically focusing on its constant and temporal components, i.e. it is defined as $P_{t}=100(\exp\{10(\overline{\delta}_{1}+\delta_{1t})\}-1)$. For each scenario, the black solid line represents the space-averaged true exposure effect. The estimate of $P_{t}$ obtained from SDGLMC is shown in red, with the solid line indicating its average, and the shaded area indicating the percentiles $2.5$ and $97.5$ of its distribution over 100 replicates.
@@
The dashed gray and dashed green lines represent the pooled estimates from the Null and GLMadj models, averaged over the 100 replicates. Whereas the Null model is evidently biased in all cases, GLMadj and SDGLMC significantly reduce the confounding bias in Scenario S1, although without entirely removing it. In Scenario S2, the confounding bias reduction is still achieved by SDGLMC compared to the Null model, but to a lesser extent. This could be because the unmeasured confounder varies at high frequencies in time, so it cannot be captured by the time-varying intercepts that only account for gradual changes. Also, the difference $X_{it}-\widehat{X}_{it}$ is approximately as rough as $Z_{it}$, and, due to their correlation, a portion of the variability of the unmeasured confounder is explained by the local exposure effect. Scenario S3 is very similar to Scenario S1, suggesting that the lack of smoothness in space of the unmeasured confounder does not seem to alter the performance (in terms of confounding bias) of the proposed approach.
@@
Figure 3: Simulation Study 1: space-averaged percent change of risk associated with a 10-unit increase in the exposure, $P_{t}$, for three scenarios and for the Null (dashed gray line), GLMadj (dashed green line), and SDGLMC (red solid line) models. The shaded area represents the percentiles $2.5$ and $97.5$ of the posterior means of $P_{t}$ obtained from SDGLMC over 100 replicated datasets. The black solid line represents the true exposure effect.
With regard to the variability of the exposure effect over space, we have computed the posterior means of the spatial effects, $\tilde{\bm{\delta}}_{1}$, for each scenario. The average of these posterior means over 100 replicated datasets are shown in Figures 46 and are compared to the (averaged) posterior means of the differences between area-specific and pooled estimates – i.e., the same quantity depicted in Figure 9 – obtained from the competing models.
All methods recover the simulated spatial pattern, with lower values in the middle of the study area. SDGLMC provides accurate estimates in all scenarios, while the Null and GLMadj models tend to underestimate the spatial variability in some cases.
Figure 4: Simulation Study 1: area-specific deviations from the time-varying percent change of risk associated with a 10-unit increase in the exposure, under Scenario S1.
Figure 5: Simulation Study 1: area-specific deviations from the time-varying percent change of risk associated with a 10-unit increase in the exposure, under Scenario S2.
Figure 6: Simulation Study 1: area-specific deviations from the time-varying percent change of risk associated with a 10-unit increase in the exposure, under Scenario S3.
The results above are confirmed also by Tables E.4–E.6 in the Supplementary Material, where the average mean squared error ($MSE_{avg}$), the maximum absolute bias ($MAB$), and the average mean absolute error ($MAE_{avg}$) are reported for each model and each scenario (see the Supplementary Material, Section E.2, for definitions). We also report these metrics for each component of the exposure effect to assess their contribution to the overall error. In general, we note that SDGLMC outperforms the competing models in all scenarios.
@@
Simulation Study 2: Recovery of the Local Exposure Effect ($f^{\prime}$) and Model Reparameterization
@@
In the second simulation study, described in Section F of the Supplementary Material, we want to assess whether the proposed approach is able to recover the first derivative, $f^{\prime}(\cdot;\bm{\gamma}_{f})$, when the functional form is assumed known. The data are generated using Equations (1)–(2), and we show that SDGLMC is able to capture the dynamics of the exposure effect, but there remains a residual bias which is due to confounding.
Another objective of this second simulation study is to motivate, from a practical viewpoint, the decomposition of the exposure effect into local-scale and large-scale components described in Equation (3). We compare SDGLMC with a reparameterized model that does not remove the overall trend from the exposure. As expected from the discussion in Section 3, the results suggest that the proposed approach is more effective in reducing the confounding bias and always has lower $MSE_{avg}$, $MAB$ and $MAE_{avg}$ values.